The sum of two numbers is 40. If two is added to the larger number,the result is equal to twice the smaller number. What are the two numbers? (simultaneous equation)

Answer by rothauserc(4718)

let x, y be the two numbers with x < y
:
1) x + y = 40
2) y + 2 = 2x
:
solve equation 1 for y
:
y = 40 - x
:
substitute for y in equation 2
:
40 - x + 2 = 2x
:
3x = 42
:
x = 14
:
y = 40 - 14 = 26
:
************************************
the two numbers are 14 and 26
:
check the answer
:
14 + 26 = 40
:
26 + 2 = 2 * 14
:
28 = 28
:
our answer checks
************************************
: